# N-DriverMotion: Driver motion learning and prediction using an event-based camera and directly trained spiking neural networks

### v1

@@ 4.3 Network configuration and training
To enable adequate training, the dataset has been converted where events would be converted to tensors consisting of x coordinate, y coordinate, polarity, and timestamp. Also, we have set the sampling time as 2.0 seconds. With such a form, it can now be conveyed as input to the CSNN during runtime.
All of the applications were implemented and mapped to GPU/CPU to be executed as the simulation for Loihi 2, with the Lava version 0.4.4 software framework using the Lava-DL library for training. Our CSNN models were trained with the event-based dataset using SLAYER provided by Lava-DL.
We have trained multiple configurations of CNN to compare the accuracy. Each network has been trained for 200 epochs with a single batch size due to the dataset’s high resolution. We have selected CUBA LIF (CUrrent BAsed Leaky Integrate and Fire) as the base neuron model for the spiking CNN models. For optimization, the learning rate has been set to 3$\times$10-3 and used ADAM for the optimizer. For the loss, we have chosen spike rate loss. The network configuration for the 4-layered CSNN and the neuron parameters are shown in Table 2.
@@ 4.4 Results
The results from training models are shown in Figure 4 and Table 3. In the benchmark dataset for evaluation of our models, we used the DVS Gesture recognition dataset, comprising 11 hand gesture categories from 29 subjects under 3 illumination effects [19]. We tested three different SNN configurations: 1) having all layers composed of dense connection, 2) containing a convolutional spiking layer with 2 fully-connected layers with the pooling kernel’s size being 8, and 3) the same model as 2 with the pooling kernel’s size being 18. For all models, we have selectively used only the first 2.0 seconds out of 3.0 seconds of motion streams for each class to classify the actions. Also, the temporal resolution was set to 5 milliseconds to increase the training efficiency. For each experiment, we have recorded the best loss and accuracy of the SNNs for the training and inference steps, and the confusion matrix for the corresponding inference steps. The confusion matrix in Figure 4 shows that the normal situation (label 3) is classified from the hazardous situations. Figure 5 illustrates the classification result with output spikes captured on the output layer classifying 13 driver motions.
@@
The system from the proposed CSNN model with 4 layers showed the highest accuracy, which reached the accuracy of 94.04% in the test. The inclusion of the pooling and convolution layer improved the accuracy performance by almost 8% over the 3-layered model only comprised of fully-connected layers. Even though our model is simplified with restrictions applied to resolve the out-of-memory problem caused by the high-resolution event data for training and testing, it was able to show comparable results. We note that our proposed dataset has more categories for classification and does not have repetitive actions, unlike the DVS Gesture recognition dataset.
5 Conclusions
In this paper, we propose N-DriverMotion for driver motion learning and prediction using a high-resolution event-based camera and directly trained spiking neural networks. As demonstrated by the experimental results above, driver motions captured by the Prophesee event-based camera could be classified with high accuracy (94.04%) using a simplified convolutional spiking neural network with low complexity. From this result, N-DriverMotion can contribute to developing safety systems for driver motion recognition in challenging conditions such as low-light environments or tunnels.

### v2

@@ 4.4 NETWORK CONFIGURATION AND TRAINING
To enable adequate training, the dataset was converted from binary format into tensors consisting of x-coordinates, y-coordinates, polarities, and timestamps. We also set the sampling time to 2.0 seconds, as the core movements were mostly within this range. More importantly, we wanted to check whether our proposed model is capable of classifying driver motions in a short amount of time, as in real-time applications, such systems should detect abnormal activities rapidly. In this form, the data can now be used as input to the CSNN during runtime.
All of the applications were implemented and mapped to GPU/CPU to be executed as the simulation for Loihi 2, using the Lava version 0.4.4 software framework and the Lava-DL library for training. Our CSNN models were developed with neuron models and trained with the event-based dataset using the SLAYER API, which were all provided by the Lava-DL.
We trained multiple CNN configurations to compare accuracy results. Each network was trained for 200 epochs with a single batch size due to the dataset’s high resolution. We selected CUBA LIF as the base neuron model for the spiking CNN models. For optimization, the learning rate was set to 3$\times$10-3 and ADAM was used as the optimizer. For the loss function, we chose spike rate loss. The network configuration for the 4-layered CSNN and the neuron parameters are shown in Table 3
@@ 5 Dense 13 N/A N/A N/A
4.5 EDGE DEVICE CONFIGURATIONS AND INFERENCE
In this section, we provide a detailed description of various AI accelerators that are specialized for use in edge devices. We utilized two different edge AI accelerators: Jetson Xaior NX, and Intel Loihi 2 to compare overall energy efficiency with respect to image processing delays.
4.5.1 Jetson Xavior NX
The Jetson Xavior NX is a fully featured development board equipped with a 6-core ARM CPU and a 48 tensor core processing unit supporting a clock frequency up to 1.1 GHz, along with 16 GB of memory. To minimize power usage, we operated on the standard module, where the GPU’s clock frequency is set to 0.67 GHz. All inference configurations were the same as with the RTX 3080, except for the library version. Due to the restriction of the Python version, inference on the Jetson Xavior NX was executed with a lower version of the lava-dl (0.4.0) library.
Figure 6: Pipeline for Inference of CSNN on the Loihi 2 processor. The blocks that are part of the main processes are colored as gray, while the green blocks are not part of the main process.
Table 4: Network configuration of a 4-layered CSNN for Loihi 2
layer map size channels kernel padding
Input 40 x 40 2 N/A N/A
Conv 40 x 40 4 5 2
Flatten 40 x 40 4 N/A N/A
Dense 512 N/A N/A N/A
Dense 13 N/A N/A N/A
4.5.2 Intel Loihi 2
A Neuromorphic chip is one of the edge AI accelerators that particularly targets the implementation of SNNs. Among various neuromorphic chips, we used Loihi 2 to observe the overall power consumption of the benchmark three fully connected layer SNN model, and the proposed CSNN model, as Loihi 2 adopts CUBA LIF as the base neuron model best fitting all of the proposed models [26]. For this experiment, we used the Oheo Gulch, one of the Loihi 2-based neuromorphic systems that contains a single-socketed Loihi 2 chip. Due to the limited availability of Loihi 2 chips, along with the nature of large input and model size, some adjustments to the input and models were made. We downsampled the input from 720x720 to 40x40 with max pooling and extracted the motion stream time from total of 3.0 seconds to 1.7 seconds. For the three fully connected layer model, there were no changes. For the proposed CSNN model, we excluded the CUBA pooling layer since the input had already been downsampled and reduced the number of features extracted from 16 to 4. The description of the network model on Loihi 2 is shown in Table 4. We were able to load the three fully connected layer model onto a single Loihi 2 chip, whereas the proposed CSNN model required a total of 4 Loihi 2 chips to be loaded.
To enable inference on Loihi 2, it is necessary to create a pipeline that can process spatio-temporal data for classification tasks on Loihi 2. We created a pipeline to conduct inference on Loihi 2, which is shown in Figure 6. The event data was stored in a buffer block named RingBuffer, where it passed the frame data one at a time to Loihi 2. As the frame data was passed to Loihi 2 the last layer in the model produced spikes that generated the classification result for the single frame. The spikes were cumulatively stored in the OutputProcess block until every frame in the event data was processed. The final classification was conducted in a Winner-Takes-All (WTA) fashion, where the class with the most spikes was the predicted value.
The total description of how the models were converted and converged at the Oheo Gulch neuromorphic hardware is shown in Table 5 and Table 6, where each column indicates the percentage of the total usage for the input axons, neuron group, neurons, synapses, axonal mapping, and axon memory, respectively.
Table 5: Resource utilization for a 3-fully connected layer model on Loihi 2
AxonIn NeuronGr Neurons Synapses AxonMap AxonMem Total Cores
3.20% 12.50% 0.32% 12.80% 0.08% 0.00% 12.94% 1
3.20% 12.50% 1.56% 73.60% 0.40% 0.00% 62.09% 8
25.31% 12.50% 0.12% 50.62% 0.03% 0.00% 60.80% 103
Total 112
Table 6: Resource utilization for the proposed CSNN model on Loihi 2. The amount of core utilization for the proposed model is more than twice of the 3-fully connected layer model due to the convolutional layer acting as the bottleneck of the model
AxonIn NeuronGr Neurons Synapses AxonMap AxonMem Total Cores
1.60% 12.50% 0.32% 6.40% 0.08% 0.00% 6.54% 1
40.00% 12.50% 0.02% 40.00% 0.01% 0.00% 64.02% 256
0.01% 12.50% 62.52% 0.11% 16.00% 0.00% 25.70% 3
Total 260
@@ 4.6.1 Accuracy on 13-class prediction
The results for 13 motion classifications from the training models are shown in Figure 7 and Table 7. In the benchmark dataset for evaluating our models, we used the DVS Gesture recognition dataset, comprising 11 hand gesture categories from 29 subjects under 3 illumination conditions [20]. We tested three different SNN configurations: 1) all layers composed of dense connection, 2) a convolutional spiking layer with 2 fully connected layers, with the pooling kernel size set to 8, and 3) the same model as 2, except with the pooling kernel size set to 18. For all models, we selectively used only the first 2.0 seconds out of 3.0 seconds of motion streams for each class to classify the actions. Additionally, the temporal resolution was set to 5 milliseconds to increase training efficiency. For each experiment, we recorded the best loss and accuracy of the SNNs for both training and inference steps, as well as the confusion matrix for the corresponding inference results. The confusion matrix in Figure 7 shows that the normal situation (label 3) is distinguished from the hazardous situations. Figure 8 illustrates the classification results with output spikes captured on the output layer, classifying 13 driver motions.
@@
The system from the proposed CSNN model with 4 layers showed the highest accuracy, reaching 94.04% in the test. The inclusion of the pooling and convolution layers improved accuracy performance by almost 8% over the 3-layered model, which was composed only of fully connected layers. Even though our model is simplified with restrictions applied to resolve the out-of-memory problem caused by the high-resolution event data for training and testing, it was able to demonstrate comparable results. We note that our proposed dataset has more categories for classification and does not include repetitive actions, unlike the DVS Gesture recognition dataset.
4.6.2 Accuracy on unexpected abnormal driver motion classification task
The results for binary classification of unexpected abnormal motions and normal motions are displayed in Figure 9. Except for 4 normal motions which were classified as abnormal driving motions, the model was highly capable of distinguishing between two categories, reaching an accuracy of 97.24%. One significant point to note is that there were no abnormal motions misclassified as normal motions, which demonstrates the robustness of our proposed CSNN model in detecting such unexpected motions.
Figure 9: Classification result for unexpected abnormal driver motion. The model detects all the abnormal actions correctly, demonstrating high sensitivity in detecting these actions, and also classifies normal driver motions with high accuracy.
4.6.3 Latency and Energy
To measure the power consumption for the Nvidia RTX 3080, we used the NVIDIA Management Library (NVML), which allows us to extract the throughput and energy consumption with respect to single event inference. Due to the fast inference time of the GPU, we calculated the latency by averaging the values, dividing the overall inference time by the total number of inference samples. Energy consumption was also measured by averaging the total energy over the total time of inference.
Nvidia Jetson Xavier NX does not support NVIDIA Management Library, unlike the RTX 3080 GPU. To resolve this issue, we utilized jetson__\__stats library, which enables us to keep track of the energy consumptions of the GPU by logging the statistics as a CSV file. Similar to the RTX 3080, the calculation for latency was done by averaging the values. The experiment was executed with no additional GPU processes other than the measurement process being loaded.
In the case of Loihi 2, the LAVA framework supports profiler tools which enables us to obtain various measurement values related to power, execution time, counts of neurocore activities, and neurocore SRAM utilization rates. To get precise latency and energy dissipation results, we ran the model 1 million times, which provided enough time for the profiler to measure the overall costs. All experiments for latency and energy comparisons between the GPUs and Loihi 2 were conducted with equal models, both adjusted for Loihi 2. The results for throughput and various energy-related measurements for single sample inference are shown in Table 8.
The overall energy consumption for Loihi 2 for 1 million inferences was 1.82 J which was highly comparable to that of both GPUs, where the RTX 3080 required over 500 J, and the Jetson Xavior NX required 1.13 J for a single inference on the CSNN model. The throughput, however, showed that while Loihi 2 processed 40.835 samples per seconds, the RTX 3080 processed 66.214 samples per seconds. We believe that this was due to the inference method of Loihi 2, which did not use matrix multiplication, but only addition for the input spike data. The Jetson Xavior NX performed very poorly on throughput, with only 5.46 samples being processed per second.
To compare the balance of low energy and fast runtimes of the devices, we calculated the energy-delay product (EDP), where the energy consumption and performance of the devices were weighted equally [42]. The calculation for EDP was conducted by multiplying the total energy by the latency for a single sample inference. In the EDP comparison between Loihi 2 and the two GPUs, we observed that Loihi 2 had a better EDP than the GPUs for both models, with the 3FCN model being 1.8 million times and the CSNN model being 20,721 times more efficient compared to the RTX 3080, and the 3FCN model being 16,278 times and the CSNN model being 541 times more efficient than the Jetson Xavior NX.
Table 8: Comparisons between GPUs and Loihi 2 on inference cost for a single sample. It is clear that despite the throughput of the RTX 3080 being lower, the energy delay product (EDP) of Loihi 2 is much lower than that of both GPUs, demonstrating the high energy efficiency of Loihi 2 for SNN models
Hardware Inference Cost Per Single Sample
Latency Throughput Energy EDP
(ms) (sample/s) Total (mJ) Dynamic(mJ) $\mu$Js
NDriverMotion 3FCN Nvidia RTX 3080 11.544 86.627 5.163 x $10^{5}$ N/A 5.94 x $10^{6}$
Nvidia Jetson Xavier NX 83.3 12.011 6.29 x $10^{2}$ N/A 5.24 x $10^{4}$
Loihi 2 Oheo Gulch 1.33 x $10^{3}$ 442.364 2.42 x $10^{-3}$ 1.89 x $10^{-4}$ 3.219
NDriverMotion CSNN Nvidia RTX 3080 15.103 66.214 5.218 x $10^{5}$ N/A 7.88 x $10^{6}$
Nvidia Jetson Xavier NX 1.83 x $10^{2}$ 5.4617 1.13 x $10^{3}$ N/A 2.06 x $10^{5}$
Loihi 2 Oheo Gulch 1.441 x $10^{4}$ 40.835 2.64 x $10^{-2}$ 7.46 x $10^{-4}$ 380.29
4.7 FUTURE WORK
While our proposed driver motion recognition system achieves significant improvements in energy efficiency and accuracy, there are several directions for future work to further enhance its performance and applicability. First, we plan to extend the N-DriverMotion dataset by increasing the diversity of driving scenarios, including more complex driving conditions, varied motion patterns, and a larger pool of participants. This will enable the model to generalize more effectively to real-world driving environments. Additionally, we plan to explore approaches for enhancing safety for pedestrians and vehicles in autonomous driving conditions. Another important area of exploration is the integration of online learning mechanisms within the neuromorphic system, allowing the model to adapt dynamically to changing driving conditions and new motion patterns in real time. To achieve this, we will explore unsupervised learning methods based on the spike-timing-dependent plasticity (STDP) learning rule. Moreover, we will evaluate the system’s scalability and performance across different hardware platforms, including more neuromorphic processors and edge devices. Lastly, we aim to optimize the neuromorphic pipeline to address challenges such as latency reduction and increased robustness in highly dynamic and unpredictable driving scenarios, further enhancing the system’s practicality in real-time autonomous driving applications.
5 CONCLUSION
In this paper, we propose N-DriverMotion, a newly collected event-based dataset designed to learn and predict driver motions using a neuromorphic vision system. Our system consists of an event-based camera that captures driver movements as spike inputs and an optimized convolutional spiking neural network (CSNN) capable of efficient inference on 720 x 720 event streams without the need for computationally expensive preprocessing. The dataset includes 13 driver motion categories, classified by direction, lighting conditions, and participants, including challenging environments like low-light conditions and tunnels.
Our experiments show that the system achieved a high accuracy of 94.04% in recognizing driver motions, even under difficult conditions such as varying light and directional inputs. This demonstrates that our proposed four-layer CSNN, designed for training and inference on energy- and resource-constrained platforms, can meet performance requirements. In addition, our experiments display the capability of distinguishing between unexpected abnormal driver motions and normal driver motions with an accuracy of 97.24%. Our system was especially able to classify all the abnormal driver motions, indicating high reliability in sensing various abnormal driver motions when deployed in real applications. We also demonstrated that our system was comparable to other applications that implemented direct use of event vision systems with SNNs, such as the DVS Gesture recognition task, where our proposed system achieved similar accuracy with a less complex model structure. When deployed on Intel’s Loihi 2 neuromorphic processor, the system demonstrated significant energy efficiency, with an energy-delay product (EDP) that was over 20,721 times more efficient than the RTX 3080 and 541 times more efficient than the Jetson Xavior NX. This level of efficiency is critical for real-time, on-device AI applications where power consumption is a key constraint.
By eliminating the need for time-consuming preprocessing and significantly reducing power consumption while maintaining high accuracy, our system can advance the safety and efficiency of autonomous driving systems using neuromorphic vision technology. Furthermore, the N-DriverMotion dataset provides a valuable resource for future research in driver behavior prediction, particularly in AI-driven systems requiring a high-resolution event dataset.